### v1

@@ III-C4 Experimental Pipelines and Labeling Approach
All pipelines were executed with two distinct labeling approaches, which we refer to as $\ell_{1}$ and $\ell_{2}$ (resulting in a total o twenty-four final fall-detection models). In both approaches, we worked with binary labels, namely, “fall” and “not-fall”, as mentioned in Section III-C3. However, $\ell_{1}$ considers the activities OM_6, OM_7, and OM_8 as fall activities. In contrast, in $\ell_{2}$, activities OM_6, OM_7, and OM_8 are considered non-falls. The resulting percentages of the positive class (“fall”) in labeling approaches $\ell_{1}$ and $\ell_{2}$ are $11.81\%$ and $9.82\%$, respectively.
III-C5 Hyperparameter search
To determine the most suitable hyperparameter configuration for the pre-processed data and the fall detection problem, and to ensure a fair comparison between the networks trained in the time and frequency domains, we used Bayesian optimization. This approach avoids empirical hyperparameter choices, ensuring that all networks in the experimental pipelines are efficiently optimized. The suggested values for the considered hyperparameters of a CNN1D architecture are presented in Table V.
TABLE V: Suggested values for considered hyperparameters.
@@ III-C5 Hyperparameter search
Furthermore, the value for the n_trials parameter of Optuna was set to 20. This parameter represents the number of iterations or experiments conducted during the search for the best hyperparameters. It is a measure of how many times the Bayesian optimization algorithm explores the search space to find the best combination of hyperparameters that maximizes the model’s MCC. After execution, the “trial” with the best hyperparameter selection (best trial) was identified for each pipeline. Subsequently, the neural networks were trained 20 times using the “best trial”, with the previously separated test sets. The results related to these trainings can be found in Section IV.
After performing Bayesian optimization for each pipeline, we obtained the hyperparameter configurations shown in Table VI (for labeling approach $\ell 1$) and Table VII (for labeling approach $\ell 2$). These hyperparameter combinations were used in the training sessions.
TABLE VI: Optimized CNN1D architectures for different pipelines in labeling approach $\ell 1$.
@@
TABLE VII: Optimized CNN1D architectures for different pipelines in labeling approach $\ell 2$.
@@
IV Results and Discussion
This section presents the conducted experiments and corresponding results. We performed experiments on the several variations of datasets (described in Section III-C). All the experiments were conducted on a server with two Nvidia GeForce GTX2080 GPUs, each one with 16GB memory. All the experiments were implemented in Python using the open-source framework TensorFlow (https://www.tensorflow.org) for model training. Bayesian hyperparameter optimization was performed with the help of the open-source Optuna framework (https://optuna.org). We begin by explaining the evaluation metrics. After that, we describe the results and a corresponding analysis.
@@ IV Results and Discussion
The best results for labeling approaches ($\ell 1$ and $\ell 2$) are provided in Table VIII and in Table IX, respectively. The number of true positives, true negatives, false positives, and false negatives obtained in the several experimental variations are also presented as separate tables in Appendix A.
TABLE VIII: Best results for labeling approach $\ell_{1}$.
@@
TABLE IX: Best results for labeling approach $\ell_{2}$.
@@
Considering the results reported in Table VIII and in Table IX, labeling approach $\ell_{2}$ consistently outperformed labeling approach $\ell_{1}$ in most pipelines, especially for the chest device. This reflects that labeling activities OM_6, OM_7, and OM_8 (corresponding to military transition activities to the lying position) as non-fall events improves the generalization capacity of the models.
@@ Appendix A Complete results
Table X and Table XI present the number of true positives, true negatives, false positives, and false negatives obtained in the several experimental variations.
TABLE X: Best results for labeling approach $\ell_{1}$.
@@
TABLE XI: Best results for labeling approach $\ell_{2}$.

### v2

@@ 3.3.4 Experimental Pipelines and Labeling Approach
All pipelines were executed with two distinct labeling approaches, which we refer to as $\ell_{1}$ and $\ell_{2}$ (resulting in a total o twenty-four final fall-detection models). In both approaches, we worked with binary labels, namely, “fall” and “not-fall”, as mentioned in Section 3.3.3. However, $\ell_{1}$ considers the activities $\text{OM}_{6}$, $\text{OM}_{7}$, and $\text{OM}_{8}$ as fall activities. In contrast, in $\ell_{2}$, activities $\text{OM}_{6}$, $\text{OM}_{7}$, and $\text{OM}_{8}$ are considered non-falls. The resulting percentages of the positive class (“fall”) in labeling approaches $\ell_{1}$ and $\ell_{2}$ are $11.81\%$ and $9.82\%$, respectively.
3.3.5 Hyperparameter search
To determine the most suitable hyperparameter configuration for the pre-processed data and the fall detection problem, and to ensure a fair comparison between the networks trained in the time and frequency domains, we used Bayesian optimization. This approach avoids empirical hyperparameter choices, ensuring that all networks in the experimental pipelines are efficiently optimized. The suggested values for the considered hyperparameters of a CNN1D architecture are presented in Table 5.
Table 5: Suggested values for considered hyperparameters.
@@ 3.3.5 Hyperparameter search
Furthermore, the value for the n_trials parameter of Optuna was set to 20. This parameter represents the number of iterations or experiments conducted during the search for the best hyperparameters. It is a measure of how many times the Bayesian optimization algorithm explores the search space to find the best combination of hyperparameters that maximizes the model’s MCC. After execution, the “trial” with the best hyperparameter selection (best trial) was identified for each pipeline. Subsequently, the neural networks were trained 20 times using the “best trial”, with the previously separated test sets. The results related to these trainings can be found in Section 4.
After performing Bayesian optimization for each pipeline, we obtained the hyperparameter configurations shown in Table 6 (for labeling approach $\ell 1$) and Table 7 (for labeling approach $\ell 2$). These hyperparameter combinations were used in the training sessions.
Table 6: Optimized CNN1D architectures for different pipelines in labeling approach $\ell 1$.
@@
Sc1AccT 37 3 4 2 107 0.2 3 0.6
Sc1AccF 303 3 3 2 64 0.2 0.0001 0.6
Sc1GyrF 83 2 2 3 62 0.5 1 0.5
Sc1GyrT 293 3 2 3 188 0.2 0.0001 0.7
Sc2AccF 82 5 4 2 99 0.2 1 0.8
Sc2AccT 100 4 4 3 89 0.1 0.0003 0.9
Sc2GyrF 566 4 3 1 294 0.2 0.0003 0.5
Sc2GyrT 256 6 2 3 137 0.4 0.0006 0.7
Sc_3_F 287 5 2 2 226 0.2 0.0003 0.7
Sc3T 8 5 3 1 255 0.2 0.0006 0.6
Sc4F 242 3 3 3 76 0.2 0.0006 0.7
Sc4T 16 3 3 2 134 0.1 3 0.8
Table 7: Optimized CNN1D architectures for different pipelines in labeling approach $\ell 2$.
@@
Sc1AccT 253 6 3 2 206 0.2 0.0006 0.8
Sc1AccF 217 5 4 3 290 0.2 0.0006 0.5
Sc1GyrF 59 3 3 2 77 0.1 3 0.5
Sc1GyrT 299 3 3 3 102 0.2 0.0006 0.8
Sc2AccF 160 4 2 2 163 0.2 0.0003 0.6
Sc2AccT 177 3 3 3 78 0.2 0.0003 0.8
Sc2GyrF 109 4 3 1 144 0.2 0.0006 0.7
Sc2GyrT 15 6 2 2 92 0.5 1 0.6
Sc3F 550 5 2 3 92 0.2 1 0.5
Sc3T 33 2 3 3 67 0.4 6 0.6
Sc4F 88 6 2 3 64 0.5 0.0003 0.5
Sc4T 9 6 2 1 74 0.1 0.01 0.9
@@
4 Results and Discussion
This section presents the conducted experiments and corresponding results. We performed experiments on the several variations of datasets (described in Section 3.3). All the experiments were conducted on a server with two Nvidia GeForce GTX2080 GPUs, each one with 16GB memory. All the experiments were implemented in Python using the open-source framework TensorFlow (https://www.tensorflow.org) for model training. Bayesian hyperparameter optimization was performed with the help of the open-source Optuna framework (https://optuna.org). We begin by explaining the evaluation metrics. After that, we describe the results and a corresponding analysis.
@@ 4 Results and Discussion
The best results for labeling approaches ($\ell 1$ and $\ell 2$) are provided in Table 8 and in Table 9, respectively. The number of true positives, true negatives, false positives, and false negatives obtained in the several experimental variations are also presented as separate tables in Appendix A.
Table 8: Best results for labeling approach $\ell_{1}$.
@@
Table 9: Best results for labeling approach $\ell_{2}$.
@@
Considering the results reported in Table 8 and in Table 9, labeling approach $\ell_{2}$ consistently outperformed labeling approach $\ell_{1}$ in most pipelines, especially for the chest device. This reflects that labeling activities $\text{OM}_{6}$, $\text{OM}_{7}$, and $\text{OM}_{8}$ (corresponding to military transition activities to the lying position) as non-fall events improves the generalization capacity of the models.
@@ Appendix A Complete results
Table 10 and Table 11 present the number of true positives, true negatives, false positives, and false negatives obtained in the several experimental variations.
Table 10: Best results for labeling approach $\ell_{1}$.
@@
Table 11: Best results for labeling approach $\ell_{2}$.